The market for tiger shrimp is initially in equilibrium at a price of $15 in a quantity of 80 ...now suppose producers decide to cut output to 40 in order to raise the price to $18. What is the value of the deadweight loss at a price of $18?
1
Expert's answer
2020-03-09T09:58:38-0400

The deadweight loss is:

DWL = 0.5×(80 - 40)×(18 - Ps) = 20×(18 - Ps), where Ps is the price from the supply curve at Q = 40 units.